OpenStudy (anonymous):

if f(x) = (x ^-x - x^0 + 2^x), then f (3) is equal to a -22 b -21 c 7 1/27 d 8 1/27

OpenStudy (anonymous):

substitute 3 for x and..

OpenStudy (anonymous):

f(x)=(x^-x-x^0+2^x) f(3) = 3^-3 - 3^0 + 2^3 f(3) = 1/27 - 1 + 8 f(3) = 7 1/27

OpenStudy (anonymous):

thankyou